OpenStudy (anonymous):

Find the volume of a cube with the given side length. 7 units A. 49 square units B. 49 cubic units C. 343 square units D. 343 cubic units

OpenStudy (anonymous):

@iGreen.

OpenStudy (nerdgirl):

Use this formula to solve it out: V = a^3

OpenStudy (anonymous):

^^^^ \(V = a^3\) \(V = 7^3\) Find the cube of 7. \(7^3 \rightarrow 7 \times 7 \times 7\)

OpenStudy (anonymous):

Recall that the volume of a cube is:\[V = a ^{3}\]

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Yes, 343 is correct.

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

theres 2 343